OpenStudy (anonymous):

Simplify the given equation. 6 - (3x + 10) + 4(2 - x) = 15

OpenStudy (anonymous):

6 - (3x + 10) + 4(2 - x) = 15 6 - 3x - 10 + 8 - 4x = 15 - 7x + 4 = 15 4 - 7x = 15 Answer: 4 - 7x = 15